OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 2x^8 + 44x^4 + 242 i already got to this part 2(x^8+22x^4+121) someome please explain to me what the next step is?

OpenStudy (anonymous):

The way i would solve this question is use the method of substitution then solve it quadratically, then refer it back to its original state

OpenStudy (anonymous):

so you need to let x^8=m^2, so x^4=m

OpenStudy (phi):

can you facator u^2 + 22u +121 ? with u= x^4

OpenStudy (anonymous):

2(m^2 +22m + 121)=0

OpenStudy (anonymous):

now 2(m+11)^2=0

OpenStudy (anonymous):

so m=-11

OpenStudy (anonymous):

therefore x^4=-11, and x^8=121

OpenStudy (anonymous):

have you done complex numbers?